OpenStudy (neverendingcycle):

Write the equation of the line containing the point (4, 6) and having a slope of 4 in slope-intercept form.

pooja195 (pooja195):

y=mx+b The slope is given y=4x+b To find b take the given coordinates and plug them into the equation 6=4(4)+b Solve for b once you get b plug it into this equation (the one we started with) y=4x+b <---plug here

OpenStudy (neverendingcycle):

6=4(4)+(-2)?

OpenStudy (neverendingcycle):

oops

OpenStudy (neverendingcycle):

6=4(4)+(-10)?

OpenStudy (neverendingcycle):

@pooja195

pooja195 (pooja195):

Not sure what you did there :/

pooja195 (pooja195):

lets do it together :-) 4 * 4=?

OpenStudy (neverendingcycle):

16

pooja195 (pooja195):

good now we have 6=16+b subtract 16 from both sides 6-16= ?

OpenStudy (neverendingcycle):

-10

pooja195 (pooja195):

Oh i see what u did now you were right my apologies good so b=-10 now plug it in y=4x-?

OpenStudy (neverendingcycle):

y=4x-10?

pooja195 (pooja195):

Perfect :)

OpenStudy (neverendingcycle):

so the answer is y=4x-10 right?